OpenStudy (anonymous):

Will give medal and fan Given the two functions below, what is the y-intercept of g(x)? f(x) = x2 g(x) = f(x –7)

OpenStudy (anonymous):

To find the y-intercept, we simply calculate the image of 0 ! We have : \[g(0)=f(0-7)=f(-7)=(-7)^2=49\] So the y-intercept of g is in the point (0,49)

OpenStudy (anonymous):

The x intercept?

OpenStudy (anonymous):

To find the x-intercept we have to solve the equation g(x)=0. We have : \[g(x)=0\iff f(x-7)=0 \\\iff (x-7)^2=0 \\\iff x-7=0\\\ x=7. \] So : the x-intercept of the g is in the point (7,0)